Journalist Report – February 24th

Sol 3 – Botanic emergency

Author : Nicolas WATTELLE

Definitely, a day without any complication on Mars, that may not exist…

Nevertheless, the day began quite well, I would even say that it was a very efficient work morning. One of the biggest challenges was setting up the antenna which we could not set the day before… The four martionauts chosen to do so were Valentine (she was wearing the Hololens during the whole outing today), Marine, Cerise and Léa. And, even if the task was not easy on paper, they appeared to manage very well: only one hour of the EVA was dedicated to achieve the tricky task.

This antenna is made of two big triangles (6 meters high), one in front of the other, directed towards France. It is called a Delta Loop, and ours was generously built by the radio club of Toulouse (REF 31), we were thus proud to see it set up in front of the Hab. We tested it, and the performances were encouraging, so that is good news for the future of our mission! A huge shout out to the four girls who made a perfect shot in EVA!

And then, the first unforeseen happened. Léa lost contact with her buddy beginning the last part of the EVA: her radio earphone and microphone had been unplugged while she was moving. So, writing on a paper to the three others, they decided to come back in the airlock, and the EVA was shortened. But this was nothing compared to what happened in the early afternoon…

We received a call from mission control. They discovered during the day that our GreenHab was compromised: white flies had contaminated the whole room so the future of all crops was at risk. The given instructions were clear: each Crew Member had to stop their activity and go to the GreenHab for a “big and fast gardening session”. We had to remove all the plants that were contaminated and assess the state of what can be salvaged and what was not.

First, we were worried for our disturbed schedule, but it turned out to be a great moment where we all had to work at the same place, at the same time. On top of this, the GreenHab atmosphere is always warm and really bright. Let’s say it, it was a pleasure doing it. But the result was quite sad: only few crops could had been saved, the GreenHab will not be usable for the next days. Currently, we do not know if we will be able to see fresh plants growing before the mission ends. The only kind we kept in our responsibility is the gravistropism experiment, which we are doing with school students back in France, which we moved to the ScienceDome.

At the end of the day, Elena the GreenHab officer had to write her last GreenHab report, and she will say goodbye to this special part of the station, that she had taken care of for the past four days…

Finally, we came back to our activities left uncompleted. And by the time I am writing these words, almost all the crew is helping to find ideas on how to cook the (big) harvest of the day: our lasts fresh vegetables of the mission.

Operations Report – February 24th

Crew 263 Operations Report 24-02-2022

SOL: 3

Name of person filing report: Mathéo Fouchet

Non-nominal systems: NA

Notes on non-nominal systems: NA

ROVERS

Spirit rover used: No

Hours: 157.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Atila and Shannon

Opportunity rover used: No

Hours: 228.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Atila and Shannon

Curiosity rover used: No

Hours: 168.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Atila and Shannon

Perseverance rover used: No

Hours: 84.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Atila and Shannon

General notes and comments: NA

Summary of Hab operations:

WATER USE: 19 gallons

Water (static tank): 422 gallons

Water (loft tank): 36 gallons

Water Meter: 01561007 units

Static to Loft Pump used: Yes

Static tank pipe heater: On

Static tank heater: On

Toilet tank emptied: No but I had to put a few gallons of water and digester in the toilets after an external intervention on them.

Summary of internet: NA

Summary of suits and radios: NA

Summary of GreenHab operations:

WATER USE: 8 gallons at 9am

Heater: Off

Supplemental light: Off

Harvest: During the afternoon, an emergency was declared and all the plants in the GreenHab had to be pulled out, due to an infestation of white flies. Therefore, all the crops were harvested by the crewmembers. The GreenHab has been shut down. There will be no GreenHab reports until further notice.

Aloe Vera 730g Peppers 12g Radish 195g Mint 50g Carrots 297g Asian Blend 150g Spring Onion 50g Tomatoes 705g Aurugula 150g Sage 1g Basil 35g Mixed greens 28g Salad greens 29g Aromatic herb (Romarin) 25g Mixed greens (different pot) 100g Parsley 10g

Summary of ScienceDome operations:

Dual split: Off

SOC: 72% at 6:45pm

Summary of RAM operations: NA

Summary of any observatory issues: NA

Summary of health and safety issues: See HSO report

Questions, concerns and requests to Mission Support: We put trashes in the rear airlock for pickup. Thanks a lot!

EVA Report – February 24th

Crew 263 EVA Report 24-02-2022

EVA # 3

Author: Léa Rouverand

Purpose of EVA: battery change for atmospheric experiment, installation of antenna and geology field campaign with LIBS

Start time: 9:00

End time: 10:40

Narrative: Rover batteries were checked just in case. The atmospheric experiment batteries were changed per the usual routine. The installation of the antenna for the HF wave propagation experiment was performed without any issues. One of the crew members lost their earplug at the beginning of the LIBS analysis so the EVA team leader decided to return to base by communicating through a notebook.

Destination: Atmospheric experience site (at intersection between Cow Dung Road and Entrance Road), around the Hab (antenna installation).

Coordinates (use UTM NAD27 CONUS): 518288 E, 4250586 N

Participants: Cerise Cuny, Léa Rouverand, Valentine Bourgeois, Marine Prunier

Road(s) and routes per MDRS Map: Entrance Road.

Mode of travel: Foot

Sol Summary – February 24th

Crew 263 Sol Summary Report 24-02-2022

Sol: 3

Summary Title: Critical Alert in GreenHab

Author’s name: Cerise Cuny

Mission Status: Nominal

Sol Activity Summary: The morning EVA went pretty well, until one of the participants lost her mic, so we headed back to the Hab. The setup of the antenna was a great success and we were soon able to listen our first signals.

Just after lunch we were contacted by Shannon for an emergency in the GreenHab: we had to act right away to get rid of a fly contamination, all hands on deck. Sadly we had to get rid of all the plants, harvesting everything we could. There probably won’t be enough time to grow any new plants before the end of our sim. On the bright side, this evening, diner is going to be made of fresh vegetables.

This emergency prevented Elena and Mathéo from doing their afternoon EVA that should be postponed to Sol 5.

Look Ahead Plan: In the next days we are going to try to catch up with the missed experiments of this afternoon.

Anomalies in work: GreenHab emergency.

Weather: Mostly sunny.

Crew Physical Status: Good.

EVA: Two EVA’s. The morning EVA went well. The afternoon EVA was cancelled.

Reports to be filed: Sol Summary, two EVA Request, EVA Report, Operations Report, Journalist Report, HSO Report.

Support Requested: We found a tool holder in the RAM which would fit perfectly our LIBS analyser for EVAs and would also allow us to carry it in a safer manner. Could we use it during our EVAs? We would put it in the RAM immediately after use.

EVA Report – February 23rd

Crew 263 EVA Report 23-02-2022

EVA # 2

Author: Cerise Cuny

Purpose of EVA: battery change for the atmospheric experiment.

Start time: 9:00 am
End time: 9:37 am

Narrative: Because of weather conditions, we only went out to change the batteries of the atmospheric experiments. We checked the rovers’ batteries, just in case. Everything went smoothly, no one slipped.

Destination: Atmospheric experience site (at the intersection between Cow Dung Road and Entrance Road).

Coordinates (use UTM NAD27 CONUS): 518288 E, 4250586 N

Participants: Cerise Cuny, Léa Rouverand, Valentine Bourgeois, Mathéo Fouchet

Road(s) and routes per MDRS Map: Entrance Road.

Mode of travel: Foot

Research Report – February 23rd

Name: Marine Prunier
Crew: 263
Date: 23 Feb 2022

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested (choose one): MLC-RCOS16
Objects to be Imaged this Evening: M81, Horsehead Nebula
Images submitted with this report: None
Problems Encountered: The observation of TOI 1555.01 was canceled in Skynet.

MUSK OBSERVATORY

Solar Features Observed: n/a
Images submitted with this report: n/a
Problems Encountered: n/a

Journalist Report – February 23rd

Sol 2 – A day on Mars’ North pole

Author: Nicolas Wattelle

Today was supposed to be a very busy day, and even if it didn’t go as planned, I will be happy to sleep well tonight.

The EVA program was completely full for the whole morning. Changing batteries for the field mill, putting back the LOAC that we took out yesterday (because it cannot stand too wet weather conditions…), installing the antenna for the HF waves propagation study, first look at Candor Chiasma in order to find a great spot to begin the geological experiment (and use LIBS for the first time!).

But it felt like we moved to Mars’ north pole during the night: when we woke up, there was snow everywhere, about 4cm around the MDRS. At first sight, this was really beautiful to see, but we changed our minds when we understood it was compromising our plans… Doing an EVA as long as we had planned would have been risky, so it was decided that the EVA would just aim to change the battery of the field mill (this was the only thing necessary, waiting another day to change the battery would have put at risk its lifetime…).

Also, the snow completely covered our solar panels. Usually, these work in pair with a generator, and they provide enough energy to endure a day at the MDRS, with a crew living and working. So, we had to cut our electrical consumption for the day, keeping only essential devices plugged. This explains the pictures of preparing for the EVA, only using the light coming through the narrow windows to see.

I think this is what trying to live on Mars would look like, everyone will have to get out of their comfort zone, living day to day with the recurrent hazards… But Crew 263 is here to live this experiment, so we accept our fate and hope for better days!

During the afternoon, we were almost all together in the upper deck of the Hab. We made rotations around different experiments, in which sometimes we are operators and sometimes, subjects. On my side, I was mainly focused on producing the first podcast episode, for curious students (and all other curious!) we collaborate with. We hope people on Earth are interested in our mission.

We tried for the first time the ultrasound experiment using a CNES protocol and a Sonoscanner device. It was really interesting, the method is very smart: with one hand, we handle the ultrasound probe, on the other a smartphone with Augmented Reality software. It shows where to place the probe and is thus like a game: we have to put 4 spheres in 4 cubes, and when it is done, we record the sequence. None of us are doctors, but thanks to this protocol we can manage to get great pictures of our organs! We are really grateful to CNES for letting us participate in the conception of this futuristic device!

On another topic, I have to confess something. We are not only 7 in this station… Since we have arrived, someone is there in the Hab upper deck, looking at us all day. He proudly stands at the window. His name is Tomarsto, and he is always there, during great moments and hard times. We hope he will stay there until the end of the simulation, and even more! Take a look at him in the picture of the day!

Operations Report – February 23rd

Crew 263 Operations Report 23-02-2022

SOL: 2

Name of person filing report: Mathéo Fouchet

Non-nominal systems: NA

Notes on non-nominal systems: NA

ROVERS

Spirit rover used: No
Hours: 157.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: handled by Atila

Opportunity rover used: No
Hours: 84.4
Beginning charge: 100%
Ending charge: 100%
Currently charging: handled by Atila

Curiosity rover used: No
Hours: 168.4Beginning charge: 100%

Ending charge: 100%
Currently charging: handled by Atila

Perseverance rover used: NoHours: 228.1
Beginning charge: 100%
Ending charge: 100%
Currently charging: handled by Atila

General notes and comments: NA

Summary of Hab operations:

WATER USE: 17 gallons
Water (static tank): 458 gallons
Water (loft tank): 38 gallons
Water Meter: 01560796 units

Static to Loft Pump used: Yes
Static tank pipe heater: On
Static tank heater: On

Toilet tank emptied: Yes

Summary of internet: NA

Summary of suits and radios: I have changed the fuse of suit #10 so that it can work correctly thanks to the advice of Scott Davis

Summary of GreenHab operations:

WATER USE: 4 gallons at 7:30 am and 4 gallons at 4:45 pm
Heater: On
Supplemental light: Off (will turn on automatically at 7:30 pm)
Harvest: NA

Summary of ScienceDome operations:

Dual split: Off
SOC: 86% at 6pm

Summary of RAM operations: NA

Summary of any observatory issues: NA

Summary of health and safety issues: See HSO report

Questions, concerns and requests to Mission Support: NA

Sol Summary – February 23rd

Crew 263 Sol Summary Report 23-02-2022

Sol: 2

Summary Title: Tiresome cognitive experiments.

Author’s name: Cerise Cuny

Mission Status: Nominal

Sol Activity Summary: This morning we woke up with a Mars polar landscape, and immediately knew that the EVA would be jeopardized because of the snow. In the end, it was decided with mission support that we would only get out to change the batteries of the atmospheric experiment. It was a brief EVA (40 minutes); the landscape was beautiful. The Crew members that stayed in the Hab performed the ultrasound experiment for the CNES. After the EVA, our Crew Engineer repaired the faulty suit and swept the corridors.

In the afternoon, we went through our experiments just as planned. Mathéo and Marine did our human factors experiment, we all were subjects and operators for the ultrasound experiment, and Nicolas and I were subjects of a cognitive experiment. This last one was particularly exhausting: we had to teleoperate a rover lying on the side (among other body positions). I was totally disoriented, I flipped the rover, and it took me 15 minutes to get it back on its wheels! Nicolas is also frustrated because he flipped the rover too: this experiment may have an impact on our mood.

Look Ahead Plan: If the weather allows it, we will set up the antenna during the EVA to be able to start communicating with Toulouse, France (simulating another mars station) and explore a predefined geological area of interest for LIBS analysis. The afternoon will be less intense than today, just a few experiments and hopefully some rest.

Anomalies in work: None to report.

Weather: Cloudy, windy.

Crew Physical Status: Good.

EVA: One performed, everything went fine.

Reports to be filed: Sol Summary, 2 EVA Requests, EVA Report, Operations Report, Journalist Report, Astronomy Report, HSO Report.

Support Requested: We are out of garbage bags, and digesters. Just in case it did not reach the right person yesterday: it would be more convenient for us to have a second working GPS for tomorrow’s EVA as we need to record the GPS information of the observed point. Currently, only one GPS is working, it would be better to have another one for redundancy.